OpenStudy (sonyalee77):

Fifteen percent of what is 80? To solve this problem, you use the basic percent equation. 0.15 x n = 80. A. true B.false

OpenStudy (jack1):

x times 0.15 = 80 so 80 / 0.15 = x x = ? so 0.15 times n = 80 = true

Directrix (directrix):

Fifteen percent is 15/100 = .15. .15n = 80 --> Let's test it. What is 80 divided by .15 ? @Sonyalee77

Directrix (directrix):

Use the Google calculator or the calculator on your neck.

OpenStudy (sonyalee77):

533.333

Directrix (directrix):

It looks as if this is going to be false.

Directrix (directrix):

But, let's see what the answer should be before we decide for sure.

OpenStudy (jack1):

...? false, why? @Directrix

Directrix (directrix):

I'm thinking and re-reading the problem. @Jack1

Directrix (directrix):

Fast is fine but accuracy is everything - Wyatt Earp

Directrix (directrix):

80/533.333 = .15000 or 15%. Okay, I am ready to say TRUE @Sonyalee77 What about you?

OpenStudy (sonyalee77):

I got the same thing

Directrix (directrix):

Okay. Right on.